What is the volume, in liters, of 4.00 moles of methane gas, CH4, at 18°C and 1.40 atm of pressure.   Please show working.
Answer #1

We Have

Number of Moles of methane gas (n) = 4.00 moles

Temperature T = 18 0C = 18 + 273.15 = 291.15 K

Pressure P = 1.40 atm

Gas Constant = R = 0.0821 L.atm/mol.K

Now We know the ideal gas equation

PV = nRT

So

Volume V = nRT/P

Lets calculate from given values

V = 4.00 mole \times 0.0821 L.atm/mol.K \times 291.15 K / 1.40 atm

= 68.30 L

So Volume =  68.30 Liter